# Daniela Rus1,3, Bryan Kian Hsiang Low1,2

### v1

@@ Abstract
We propose Tetris, a novel method that optimizes the total throughput of batch speculative decoding in multi-request settings. Unlike existing methods that optimize for a single request or a group of requests as a whole, Tetris actively selects the most promising draft tokens (for every request in a batch) to be accepted when verified in parallel, resulting in fewer rejected tokens and hence less wasted computing resources. Such an effective resource utilization to achieve fast inference in large language models (LLMs) is especially important to service providers with limited inference capacity. Compared to baseline speculative decoding, Tetris yields a consistently higher acceptance rate and more effective utilization of the limited inference capacity. We show theoretically and empirically that Tetris outperforms baseline speculative decoding and existing methods that dynamically select draft tokens, leading to a more efficient batch inference in LLMs.
00footnotetext: * Equal contribution.
@@
Zhaoxuan Wu*1, Zijian Zhou*1,2, Arun Verma1, Alok Prakash1, Daniela Rus1,3, Bryan Kian Hsiang Low1,2 1Singapore-MIT Alliance for Research and Technology, Republic of Singapore 2Dept. of Computer Science, National University of Singapore, Republic of Singapore 3CSAIL, Massachusetts Institute of Technology, USA
@@ 1 Introduction
Transformer-based large language models (LLMs) have shown remarkable abilities to solve different tasks across various domains, such as natural language (Zhao et al., 2023), computer vision (Yin et al., 2024), robotics (Zeng et al., 2023), code generation (rozière2024codellamaopenfoundation), among others (Maslej et al., 2024). However, the autoregressive nature of LLMs (i.e., generating one token at a time) leads to an increasingly sluggish inference speed as the model size increases.
@@
Such batch processing of user requests entails a fundamentally different optimization objective for SD compared to handling individual requests. For SD of a single request, supposing a fast draft model with negligible runtime, the objective is to maximize the draft window size as long as the target model can verify all draft tokens in parallel by fully utilizing the inference capacity. It can be naively extended to batch processing by widening the draft window for all requests until the inference capacity is reached. This is inefficient as each request may require a different optimal draft token selection due to varying difficulty in speculation (i.e., generating tokens to match the target model’s output).
@@ 2 Related Work
By employing a draft-then-verify strategy for lossless accelerations of LLM inference, SD has attracted significant attention recently (Ryu and Kim, 2024; Xia et al., 2024). Recent advancements based on SD have focused on developing more efficient draft models by producing multiple drafts for the next few tokens (Cai et al., 2024; Cheng et al., 2024; Li et al., 2024b). Additionally, some methods have optimized the speculation accuracy by aligning the draft model with the target model (Liu et al., 2024e; Zhou et al., 2024) or leveraging the target model itself to draft via techniques like layer skipping (Zhang et al., 2024). To facilitate more efficient verification, tree attention has been proposed for speedy tree-structured candidate verification (Miao et al., 2024; Spector and Re, 2023). In contrast, our work explores a complementary approach that intervenes between the draft and target models, performing strategic draft token selection to improve throughput over batched requests. Our method can be seamlessly integrated with the above techniques for a more efficient SD system.
@@ 4.2 Analysis
The proof is delayed to Section A.3. Overall, we established both per-step and global optimality of Tetris under theoretical assumptions. In practice, the drafting time can be hidden with appropriately designed pipeline (Liu et al., 2024c; Wang et al., 2024b) which parallelizes the execution of the draft model and the target model. [^1] Although, they have yet been integrated in popular battle-tested model serving frameworks such as vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2024) as of this writing. The true acceptance rates are inaccessible in practice, we thus rely on surrogate measures and show their empirical effectiveness, which we will discuss next.
@@ 4.3 Practical Implementations
The acceptance rate of a draft token depends on $\max(p_{\mathcal{M}}(d_{i})/p_{\mathcal{S}}(d_{i}),1)$. However, the Tetris manager does not have access to $p_{\mathcal{M}}(\cdot)$ before verification. In practice, we use the draft model’s output probability as a surrogate of the token acceptance rate (Kim et al., 2023; Zhang et al., 2024). We show in Section 5 that this surrogate empirically results in strong performance. Additionally, while we theoretically show that Algorithm 1 achieves a time complexity of ${\mathcal{O}}(C\log N)$ (see Section A.2), we can additionally leverage the parallelism of GPU to achieve empirical negligible overhead of using Tetris ($<0.3\text{ms}$ compared to the average draft time per token of $>2.5\text{ms}$) via the scatter_max operation directly implemented on GPU. Lastly, the autoregressive token drafting can also be parallelized across requests. Hence, drafting a batch of requests with a common window size of $k$ tokens takes the same time as a single request in practice.
@@ 5 Experiments
We evaluate the effectiveness and efficiency of Tetris against baseline methods. We first validate the necessity of dynamic draft token selection and improvement of token acceptance with Tetris in Sections 5.1 and 5.2. Then, we show the empirical end-to-end speedup in Section 5.3. We also discuss the potential further improvement in empirical results with the future implementation of speculative decoding pipelines in Section 5.4.
Settings.
We perform experiments on target models of various parameter sizes, including Vicuna-33B-v1.3, Llama-3.1-70B-Instruct, and Llama-3.1-405B-Instruct. We use Vicuna-68M and Llama-3.2-1B-Instruct as their respective draft models. Depending on the size of the models, different server configurations and tensor parallel sizes are adopted, detailed in Table 1. Tetris is evaluated for generation of answer completion for questions extracted from ShareGPT (Anon, 2023), Chatbot Arena (Zheng et al., 2023), Domain Tough Questions (YAV-AI, 2024), and synthetic tasks generated from Shakespeare’s The Sonnet. The standard speculative decoding (SD) (Leviathan et al., 2023) and dynamic speculative decoding (DSD) (Liu et al., 2024d) are baseline methods that we compare to. We vary the drafting window sizes, allowing up to 3 extra draft tokens for Tetris while keeping the same number of tokens sent for verification by the target model for fair comparison. Tetris is implemented in vLLM (Kwon et al., 2023).
@@ 5.1 Variations in Draft Quality
Figure 2: The distribution of the number of accepted tokens per speculative decoding step.
@@ 5.2 Effect of Extra Draft Tokens
which measures the quality of the draft tokens selected by Tetris. We show in Fig. 3 that increasing the number of extra draft tokens consistently increases the VSR metric across all settings. This finding confirms the effectiveness of Tetris’s strategy for draft token selection utilizing extra draft tokens. It also validates the empirical usefulness of the draft model’s output probabilities as a surrogate of the selection criteria, as stated in Section 4.3.
Figure 3: Change in VSR as the number of extra draft tokens increases. Base draft length $k=4$.
@@ 5.3 Evaluation of Tetris
We also measure the end-to-end latency of each request. This metric measures the average latency of the speculative decoding system in finishing completions, which can affect user satisfaction. We summarize the results in Table 2 and defer the figures to Section C.2. Overall, Tetris achieves up to 6.13% improvement in latency as compared to the best baseline and up to 9.32% improvement against standard SD.
@@ 5.5 Ablation Study
We also include ablations on Tetris’s improvement in verification success rate (VFS) in Section C.5, and the effect of batch size on the performance in Section C.6.
@@ 6 Conclusion and Future Work
In this paper, we study the problem of optimizing batch speculative decoding to maximize throughput in multi-request settings, such as those faced by model service providers. To this end, we propose Tetris, a novel method that efficiently selects optimal draft tokens for the LLM verification in log-linear time. We have theoretically shown that, in the absence of drafting time, Tetris achieves optimal throughput both at each decoding step and globally under reasonable assumptions about token acceptance rates. Our empirical results further validate that Tetris consistently outperforms standard speculative decoding and existing dynamic draft window selection methods, even when accounting for the extra time required for drafting extra tokens. These results highlight the potential of Tetris to improve inference efficiency in real-world model service deployments. A key future direction is adapting Tetris to tree decoding, a notable feature in recent advancements in speculative decoding (Cai et al., 2024; Li et al., 2024a, b).
@@ A.2 Running Time of Tetris
Appendix B Additional Related Work
@@ C.1 Dataset License
C.2 Plots for End-to-end Latency
@@
C.3 Plots for Projected Improvement based on TER
@@
C.4 Extension to Medusa
@@
C.5 Improvement in Verification Success Rate
@@
C.6 The Effect of Batch Size on Tetris Performance

### v2

@@ Abstract
We propose Tetris, a novel method that optimizes the total throughput of batch speculative decoding in multi-request settings. Unlike existing methods that optimize for a single request or a group of requests as a whole, Tetris actively selects the most promising draft tokens (for every request in a batch) to be accepted when verified in parallel, resulting in fewer rejected tokens and hence less wasted computing resources. Such an effective resource utilization to achieve fast inference in large language models (LLMs) is especially important to service providers with limited inference capacity. Compared to baseline speculative decoding, Tetris yields a consistently higher acceptance rate and more effective utilization of the limited inference capacity. We show theoretically and empirically that Tetris outperforms baseline speculative decoding and existing methods that dynamically select draft tokens, leading to a more efficient batch inference in LLMs. 00footnotetext: * Equal contribution.
@@
Zhaoxuan Wu*1, Zijian Zhou*1,2, Arun Verma1, Alok Prakash1, Daniela Rus1,3, Bryan Kian Hsiang Low1,2 1Singapore-MIT Alliance for Research and Technology, Republic of Singapore 2Dept. of Computer Science, National University of Singapore, Republic of Singapore 3CSAIL, Massachusetts Institute of Technology, USA Correspondence: lowkh@comp.nus.edu.sg
@@ 1 Introduction
Transformer-based large language models (LLMs) have shown remarkable abilities to solve different tasks across various domains, such as natural language (Zhao et al., 2023), computer vision (Yin et al., 2024), robotics (Zeng et al., 2023), code generation (Rozière et al., 2024), among others (Maslej et al., 2024). However, the autoregressive nature of LLMs (i.e., generating one token at a time) leads to an increasingly sluggish inference speed as the model size increases.
@@
Such batch processing of user requests entails a fundamentally different optimization objective for SD compared to handling individual requests. For SD of a single request, supposing a fast draft model with negligible runtime, the objective is to maximize the draft window size as long as the target model can verify all draft tokens in parallel by fully utilizing the inference capacity. It can be naively extended to batch processing by widening the draft window for all requests until the inference capacity is reached. This is inefficient as each request may require a different optimal draft token selection due to varying difficulty in speculation (i.e., generating tokens to match the target model’s outputs).
@@ 2 Related Work
By employing a draft-then-verify strategy for lossless accelerations of LLM inference, SD has attracted significant attention recently (Ryu and Kim, 2024; Xia et al., 2024). Recent advancements based on SD have focused on developing more efficient draft models by producing multiple drafts for the next few tokens (Cai et al., 2024; Cheng et al., 2024; Li et al., 2024b). Additionally, some methods have optimized the speculation accuracy by aligning the draft model with the target model (Liu et al., 2024e; Zhou et al., 2024a) or leveraging the target model itself to draft via techniques like layer skipping (Zhang et al., 2024). To facilitate more efficient verification, tree attention has been proposed for speedy tree-structured candidate verification (Miao et al., 2024; Spector and Re, 2023). In contrast, our work explores a complementary approach that intervenes between the draft and target models, performing strategic draft token selection to improve throughput over batched requests. Our method can be seamlessly integrated with the above techniques for a more efficient SD system.
@@ 4.2 Analysis
The proof is delayed to Section A.3. Overall, we established both per-step and global optimality of Tetris under theoretical assumptions. Similar assumptions are commonly made in the literature (Leviathan et al., 2023; Liu et al., 2024d) to enable theoretical insights (more details in Section B.3), and Tetris demonstrates strong empirical performance even when this assumption is violated, as we show later in Section 5. In practice, the drafting time can be hidden with appropriately designed pipeline (Liu et al., 2024c; Wang et al., 2024b) which parallelizes the execution of the draft model and the target model. [^1] Although, they have yet been integrated in popular battle-tested model serving frameworks such as vLLM (Kwon et al., 2023) and SGLang (Zheng et al., 2024) as of this writing. The true acceptance rates are inaccessible in practice, we thus rely on surrogate measures and show their empirical effectiveness, which we will discuss next.
@@ 4.3 Practical Implementations
The acceptance rate of a draft token depends on $\max(p_{\mathcal{M}}(d_{i})/p_{\mathcal{S}}(d_{i}),1)$. However, the Tetris manager does not have access to $p_{\mathcal{M}}(\cdot)$ before verification. In practice, we use the draft model’s output probability as a surrogate of the token acceptance rate (Kim et al., 2023; Zhang et al., 2024). We show in Section 5 that this surrogate empirically results in strong performance. While prior works such as EAGLE-2 (Li et al., 2024b) and MDSD (Hu et al., 2025) have adopted greedy token selection based on draft model probabilities on a single request, Tetris’ greedy algorithm operates at the batch level to optimize resource utilization across multiple requests, where we defer a more detailed discussion to Section B.2. Additionally, while we theoretically show that Algorithm 1 achieves a time complexity of ${\mathcal{O}}(C\log N)$ (see Section A.2), we can additionally leverage the parallelism of GPU to achieve empirical negligible overhead of using Tetris ($<0.3\text{ms}$ compared to the average draft time per token of $>2.5\text{ms}$) via the scatter_max operation directly implemented on GPU. Lastly, the autoregressive token drafting can also be parallelized across requests. Hence, drafting a batch of requests with a common window size of $k$ tokens takes the same time as a single request in practice.
@@ 5 Experiments
We evaluate the effectiveness and efficiency of Tetris against baseline methods. We first validate the necessity of dynamic draft token selection and improvement of token acceptance with Tetris in Sections 5.1 and 5.2. Then, we show the empirical end-to-end speedup in Section 5.3. We also discuss the potential further improvement in empirical results with the future implementation of speculative decoding pipelines in Section 5.4. Our code is available at https://github.com/ZhaoxuanWu/Tetris.
@@ 3 Llama-1B-FP8 (1) Llama-405B-FP8 (8) 8$\times$×H100 (640G)
Figure 2: The distribution of the number of accepted tokens per speculative decoding step for various tasks.
Figure 3: Change in VSR as the number of extra draft tokens increases. Base draft length $k$ is set to 4, results for other $k$’s are in Section C.2.
Settings.
We perform experiments on target models of various parameter sizes, including Vicuna-33B-v1.3, Llama-3.1-70B-Instruct, and Llama-3.1-405B-Instruct. We use Vicuna-68M and Llama-3.2-1B-Instruct as their respective draft models. Depending on the size of the models, different server configurations and tensor parallel sizes are adopted, detailed in Table 1. Tetris is evaluated for generation of answer completion for questions extracted from ShareGPT (Anon, 2023), Chatbot Arena (Zheng et al., 2023), Domain Tough Questions (YAV-AI, 2024), and synthetic tasks generated from Shakespeare’s The Sonnet. The standard speculative decoding (SD) (Leviathan et al., 2023) and dynamic speculative decoding (DSD) (Liu et al., 2024d) are baseline methods that we compare to. We vary the drafting window sizes, allowing up to 3 extra draft tokens for Tetris while keeping the same number of tokens sent for verification by the target model (i.e., fixing the inference capacity) for fair comparison to baseline methods. Tetris is implemented in vLLM (Kwon et al., 2023).
@@ 5.2 Effect of Extra Draft Tokens
which measures the quality of the draft tokens selected by Tetris. Fixing the total number of tokens sent for verification, we show in Fig. 3 that increasing the number of extra draft tokens consistently increases the VSR metric across all settings. This finding confirms the effectiveness of Tetris’s strategy for draft token selection utilizing extra draft tokens. It also validates the empirical usefulness of the draft model’s output probabilities as a surrogate of the selection criteria, as stated in Section 4.3.
@@ 5.3 Evaluation of Tetris
We also measure the end-to-end latency of each request, defined as the time from sending the request to receiving the final response from the vLLM server on the client side. This metric measures the average latency of the speculative decoding system in finishing completions, which can affect user satisfaction. We summarize the results in Table 2 and defer the figures to Section C.3. Overall, Tetris achieves up to 6.13% improvement in latency as compared to the best baseline and up to 9.32% improvement against standard SD.
@@ 5.5 Ablation Study
We also include ablations on Tetris’s improvement in verification success rate (VSR) in Section C.6, and the effect of batch size on the performance in Section C.7.
@@ 6 Conclusion and Future Work
In this paper, we study the problem of optimizing batch speculative decoding to maximize throughput in multi-request settings, such as those faced by model service providers. To this end, we propose Tetris, a novel method that efficiently selects optimal draft tokens for the LLM verification in log-linear time. We have theoretically shown that, in the absence of drafting time, Tetris achieves optimal throughput both at each decoding step and globally under reasonable assumptions about token acceptance rates. Our empirical results further validate that Tetris consistently outperforms standard speculative decoding and existing dynamic draft window selection methods, even when accounting for the extra time required for drafting extra tokens. These results highlight the potential of Tetris to improve inference efficiency in real-world model service deployments. A key future direction is adapting Tetris to tree decoding, a notable feature in recent advancements in speculative decoding (Cai et al., 2024; Li et al., 2024a, b). Another interesting direction to explore is to design better draft token selection techniques. Insights can be derived from existing works on data selection (Lin et al., 2024; Zhou et al., 2023; Xu et al., 2024; Zhou et al., 2024b) that find important data points. Similarly, further research can investigate how to leverage the probability distribution between the draft model and the target model to improve the selection efficiency.
@@ 7 Limitations
Acknowledgments
This research is supported by the National Research Foundation (NRF), Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. The Mens, Manus, and Machina (M3S) is an interdisciplinary research group (IRG) of the Singapore MIT Alliance for Research and Technology (SMART) centre.
@@ A.2 Running Time of Tetris
Appendix B Additional Related Work and Discussion
@@ B.1 Acceptance Rate
B.2 Greedy Algorithms for Speculative Decoding
Several existing approaches in speculative decoding employ greedy algorithms to improve efficiency, though they operate at different levels and with distinct objectives. EAGLE-2 (Li et al., 2024a) and MDSD (Hu et al., 2025), for instance, utilize greedy selection strategies within single-request, multi-draft settings. EAGLE-2 achieves speedups by generating multiple branches in a draft tree, which the top candidates are then greedily selected for verification. It requires substantial GPU resources for each request and may not translate to improved throughput in batch inference due to limited total GPU capacity. MDSD, on the other hand, focuses on increasing the acceptance rate in multi-draft speculative decoding through greedy draft sampling, an objective orthogonal to our work. Our Tetris distinguishes itself by applying a greedy algorithm at the batch level, optimizing GPU resource utilization across multiple requests rather than within a single draft tree. Due to the simplicity of greedy approaches in implementation, our method can be readily adapted to major inference frameworks (e.g., vLLM (Kwon et al., 2023)) with minimal empirical overhead. Our paper also demonstrates that Tetris is complementary to existing speculative decoding frameworks, such as Medusa (Cai et al., 2024) (see Section 5.5 and Section C.5), by introducing a novel axis of optimization not explored by prior works.
B.3 Assumptions for Analysis
Our theoretical analysis of the global optimality of Tetris in 2 is conditioned on 1. Similar assumptions are commonly made in the literature (Leviathan et al., 2023; Liu et al., 2024d) to analyze the length of generated outputs. Specifically, the pioneering work of Leviathan et al. (2023) assumes a constant acceptance rate $\alpha$ and applies a capped geometric distribution to derive the expected number of generated tokens, under a single-request setting. For a single request, this assumption is equivalent to our 1. A natural extension of the constant acceptance rate assumption to the batch inference setting would be assuming the identical acceptance rate across all requests. However, we highlight that our 1 is weaker than assuming the same constant acceptance rate across all requests, as it allows the acceptance rate to vary across requests. Moreover, even when 1 is violated, our method still demonstrates strong empirical performance as shown in Section 5 of the main paper.
Refer to caption Refer to caption Refer to caption
(a) Base $k=3$ (b) Base $k=2$ (c) Base $k=1$
Figure 7: Additional plots for the change in VSR as the number of extra draft tokens increases, supplementary to Fig. 3 where base draft length $k=4$. Here, we provide results for $k=1,2,3$.
@@ C.1 Dataset License
C.2 Additional Plots for Effect of Extra Draft Tokens
Fig. 3 demonstrates the benefit of adding extra draft tokens (from 0 to 9 extra tokens). The base draft length $k$ serves only as a reference baseline and its value does not significantly impact the observed trends in Fig. 3. We set $k=4$ based on established reasonable ranges used in existing studies (Leviathan et al., 2023; Zhang et al., 2024; Wang et al., 2024b). This choice balances the trade-off between extra compute required and additional throughput gained.
For completeness, we conducted additional experiments with $k=1,2,3$ and present the results in Fig. 7. The observed trends are consistent across different values of $k$: VSR generally increases with extra tokens, demonstrating the effectiveness of having extra draft tokens in our Tetris.
@@ C.2 Additional Plots for Effect of Extra Draft Tokens
C.3 Plots for End-to-end Latency
@@
C.4 Plots for Projected Improvement based on TER
C.5 Extension to Medusa
@@
C.6 Improvement in Verification Success Rate
@@
C.7 The Effect of Batch Size on Tetris Performance
@@
Appendix D Broader Impacts
While this research work is primarily foundational, focusing on computational performance, the resulting increase in inference speed and efficiency of Large Language Models (LLMs) could indirectly contribute to certain societal risks associated with LLMs. Making LLM inference faster and cheaper lowers the barrier to deploying these models at scale. Consequently, this could potentially accelerate the proliferation of LLM-generated content, increasing the risks of misuse such as the large-scale generation of disinformation, spam, or fake online personas, if the underlying models are deployed without adequate safeguards.
Mitigation strategies depend on responsible deployment. Developers using Tetris should employ robust safety measures, bias detection, and content filtering for the LLMs being served. Importantly, the efficiency gains from Tetris could be leveraged positively to make computational overhead for safety checks, alignment techniques, or bias mitigation measures more feasible during inference.